Delia has 115 books. If she has 4 times as many mystery books as science books, how many mystery books does she have?
Answer by checkley77(12844)

x+4x=115
5x=115
x=115/5
x=23 the number of science books.
4*23=92 mystery biooks.
Proof:
23+92=115
115=115